40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9712     Entry time: Mon Mar 10 21:16:56 2014
Author: Koji 
Type: Summary 
Category: LSC 
Subject: Composite Error Signal for ARms (5)  

After making the CDS modification, CESAR was tested with ALS.

First of all, we run CESAR with threshold of 10. This means that the error signal always used ALS.
The ALS was scanned over the resonance. The plot of the scan can be found in EricQ's elog.
At each point of the scan, the arm stability is limited by the ALS.

Using this scan data, we could adjust the gains for the TR and PDH signals. Once the gains were adjusted
the threshold was lowered to 0.25. This activates dynamic signal blending.

ALS was stabilized with XARM FM1/2/3/5/6/7/9. The resonance was scanned. No glitch was observed.
This is some level of success already.

Next step was to fully hand off the control to PDH. But this was not successfull. Everytime the gain for the TR was
reduced to zero, the lock was lost. When the TR is removed from the control, the raw PDH signal is used fot the control
without normalization. Without turning on FM4, we lose 60dB of DC gain. Therefore the residual motion may have been
too big for the linear range of the PDH signal. This could be mitigated by the normalization of the PDH signal by the TR.

ELOG V3.1.3-